OpenStudy (amber15):

Simplify the trigonometric expression. 1/1+ sin theta + 1/1-sin theta A, 2 cos^2 theta B. 2 sec^2 theta C. 2 csc^2 theta D. 2 cot^2 theta A,2 cos^2 theta B. 2 sec ^2 theta C. 2 csc^2 theta D, 2 cot^2 theta my answer is B can some one tell me if im correct i know C and A are not correct answers

OpenStudy (calculusfunctions):

\[\frac{ 1 }{ 1+\sin \theta }+\frac{ 1 }{ 1-\sin \theta }\] The LCD (least common multiple of the denominators) is (1 + sin θ)(1 − sin θ). Hence \[\frac{ 1 }{ 1+\sin \theta }+\frac{ 1 }{ 1-\sin \theta }\] \[=\frac{ 1-\sin \theta }{(1+\sin \theta)(1-\sin \theta) }+\frac{ 1+\sin \theta }{ (1+\sin \theta)(1-\sin \theta) }\] \[=\frac{ 1-\sin \theta +1+\sin \theta }{ (1+\sin \theta)(1-\sin \theta) }\] \[=\frac{ 2 }{ 1-\sin ^{2}\theta }\] \[=\frac{ 2 }{ \cos ^{2}\theta }\] \[=2\sec ^{2}\theta \] I hope that helps!

OpenStudy (amber15):

thank you

OpenStudy (amber15):

so the answer is D

OpenStudy (calculusfunctions):

You mean B, don't you?

OpenStudy (amber15):

OOPS yes

OpenStudy (calculusfunctions):

Good! Don't I get a medal? lol

OpenStudy (amber15):

yes you get a medal and thank you

OpenStudy (calculusfunctions):

Haha! I was just trying to be funny but thank you!

OpenStudy (amber15):

you did good on explaining tho

OpenStudy (calculusfunctions):

Thank you!

OpenStudy (rosedewittbukater):

@amber15 do you remember if this answer was correct?

OpenStudy (amber15):

yes it is correct

OpenStudy (rosedewittbukater):

Thank you so much!

OpenStudy (amber15):

your welcome